# Dissecting the NVIDIA Hopper Architecture through Microbenchmarking and Multiple Level Analysis

### v1

@@
Abstract
Modern GPUs, with their specialized hardware like tensor cores, are essential for demanding AI and deep learning applications. This study presents a comprehensive, multi-level microbenchmarking analysis of the NVIDIA Hopper GPU architecture, delving into its performance characteristics and novel features. We benchmark Hopper’s memory subsystem latency and throughput, comparing its L2 partitioned cache behavior and global memory access patterns against recent GPU generations, Ampere and Ada Lovelace. Our analysis reveals significant performance differences and architectural improvements in Hopper. A core contribution of this work is a detailed evaluation of Hopper’s fourth-generation tensor cores, including their FP8 precision support and the novel asynchronous wgmma instructions, assessing their impact on matrix multiply-accumulate operations. We further investigate the performance implications of other key Hopper innovations: DPX instructions for accelerating dynamic programming algorithms, distributed shared memory (DSM) for inter-SM communication, and the Tensor Memory Accelerator (TMA) for asynchronous data movement. This multi-level approach encompasses instruction-level microbenchmarks, library-level analysis of the Transformer Engine, and application-level benchmarks of tensor core performance within large language models. Our findings provide valuable, in-depth insights for software developers seeking to optimize performance and develop accurate performance models for the Hopper architecture, ultimately contributing to a deeper understanding of its potential for accelerating AI and other computationally intensive workloads.
Index Terms:
@@ Abstract
Instruction Latency, Tensor Core, PTX, Hopper, DPX, Tensor Memory Accelerator, Distributed Shared Memory
1 Introduction
@@
The increasing prevalence of GPUs has significantly accelerated a diverse range of computational workloads, from scientific computing to artificial intelligence, particularly with the rise of large language models (LLMs) such as GPT-3, with over 150 billion parameters [1]. Modern GPU architectures such as Ampere, Ada, and Hopper, incorporate specialized hardware like tensor cores and high-bandwidth memory systems, purpose-built for AI applications. As a result, GPUs have become indispensable components in high-performance computing clusters.
@@
NVIDIA maintains a two-year release cycle for new GPU architectures, ensuring the continuous integration of novel architectural features. However, the scarcity of detailed micro-architectural specifications hinders precise performance analysis, necessitating deeper investigations to understand the impact of these advancements on application performance. The evolution of tensor cores (TCs) exemplifies this trend. Introduced in the Volta architecture to accelerate deep neural networks using FP16/FP32 operations, TCs have been progressively enhanced. Ampere and Hopper continue to increase support for different precisions, providing more space for research on LLM. Thus, the reliance on assembly analysis and microbenchmarks from older architectures (Ampere and Turing) highlights the need for Hopper-specific TC research.
@@
Beyond enhanced TCs, the Hopper architecture introduces several key innovations. Dynamic Programming X (DPX) instructions accelerate dynamic programming algorithms, which often rely on comparisons (min/max operations) between previously computed results. Distributed shared memory (DSM) enables direct communication between Streaming Multiprocessors (SMs), facilitating loads, stores, and atomic operations across their respective shared memory blocks. Finally, the Tensor Memory Accelerator (TMA) provides an asynchronous data movement mechanism between hardware components. However, the implementation details and performance characteristics of the TMA remain largely unexplored in the existing literature. A comprehensive understanding of these aspects is crucial for developers seeking to fully exploit the potential of modern GPUs and optimize their AI applications. Further research is needed to fully understand the interplay of these features and their combined effect on diverse workloads.
@@
This article extends our previous conference paper [2], which explored the basic performance of the memory subsystem, a multi-level comparison of fourth-generation tensor cores with previous architectures, and a simple analysis of DPX and Distributed Shared Memory. In this work, we expand both the depth and breadth of our research. In terms of depth, we further analyze the memory subsystem. After various comparisons, we find that some traditional benchmarks are not suitable for Ampere and Hopper. We conduct new tests on the L2 partitioned cache for Ampere and Hopper, overturning previous data and conclusions. Additionally, we include energy consumption tests for wgmma in the tensor core, revealing the power impact of new instructions. We also perform a comprehensive evaluation of distributed shared memory (DSM), assessing DSM interface costs, expanding cluster numbers in latency tests, and testing throughput with different access patterns and thread block scheduling policies. Furthermore, we incorporate application-level testing to further demonstrate the advantages and limitations of DPX instructions. In terms of breadth, with the rise of asynchronous programming, we benchmark the TMA on Hopper to characterize its performance within asynchronous programming paradigms.
@@
$\bullet$ We systematically benchmark the latency and throughput of Hopper GPUs, including the memory subsystem, tensor cores, distributed shared memory (DSM), tensor memory accelerator (TMA), and dynamic programming accelerators (DPX). We compare the latency and throughput of these Hopper-specific units with their predecessors in Ampere and Ada Lovelace architectures, highlighting the performance improvements and trade-offs introduced in Hopper.
@@
2 Related Work
TABLE I: The Comparison of the studies of GPGPU Microbenchmark
Study Instruction-level Memory subsystem Energy
consumption
Tensor core Application evaluation Architectural features Target architecture
[3, 4, 5] Volta
[6, 7, 8] Turing,Volta
[9, 10, 11, 12] Asynchronous Copy Ampere,Turing
Ours DSM, TMA, DPX Ampere, Ada, Hopper
@@ 2 Related Work
The rapid evolution of GPU architectures necessitates continuous investigation into their microarchitectural details, particularly at the instruction level. This understanding is fundamental for accurate performance and power modeling [13, 14, 15, 16, 17, 18, 19, 20], the development of robust simulators [21, 22, 23], and the effective optimization of applications [22, 24, 25]. Prior research has addressed these needs across different GPU generations, but the introduction of novel features in recent architectures, such as Hopper, demands a renewed focus.
@@
Early efforts to dissect GPU microarchitecture primarily focused on older generations like Fermi, Kepler, and Maxwell [26, 27, 28]. These studies, along with investigations into performance and power characteristics [3, 6], laid the groundwork for understanding fundamental GPU behavior and informed subsequent investigations into more recent architectures like Volta and Turing. The emergence of tensor cores (TCs) in Volta, specialized hardware units designed to accelerate matrix operations crucial for AI workloads, marked a significant shift in GPU architecture and spurred a new wave of research.
@@
The importance of fused matrix multiplication accumulation (MMA) in AI has driven extensive research into TCs. Initial studies on Volta’s TCs centered on the legacy wmma API and employed benchmarks using established libraries like CUBLAS and CUTLASS [4, 5]. Further work extended this analysis to Turing, incorporating preliminary assembly code analysis of wmma instructions [3, 6]. However, these initial explorations lacked the comprehensive instruction-level microbenchmarks needed to fully characterize TC performance and numerical behavior.
@@
Subsequent research delved deeper into TC analysis on Volta and Turing, utilizing assembly-level (SASS) benchmarking to uncover optimization opportunities for matrix multiplication, particularly in half-precision [8, 25, 7]. The numerical intricacies of TCs, including rounding modes and subnormal number handling for various data types like TF32, BF16, and FP16, were also investigated [9]. However, the limitations of the wmma API, particularly its inability to fully leverage new features like sparse matrix multiplication introduced in Ampere and later architectures, became increasingly apparent [4, 9].
@@
The introduction of the mma API with Turing addressed some of these limitations and evolved to support sparse matrix multiplication (mma.sp) on Ampere and beyond. Sun et al. [12] conducted a thorough analysis of the mma API on Turing and Ampere, exploring TC performance, numerical behavior, and sparse matrix multiplication capabilities. However, the arrival of Hopper introduced further complexity with the addition of wgmma and an expanded set of mma instructions supporting a broader range of precisions. This necessitates new research to understand the performance characteristics and optimal utilization of these new instructions, a gap addressed by our work.
@@
Beyond TC-specific research, previous work has also addressed energy efficiency in GPUs using static [29, 30, 31, 32] and dynamic [16, 33, 34] analysis techniques, as well as hybrid approaches [35]. Application-level benchmarks have explored the impact of DVFS on deep learning workloads [36, 37]. Table I summarizes the existing microbenchmark studies across different GPU architectures and compares with our work. In contrast to previous research, our study provides a holistic analysis of Hopper, including its novel TCs, DPX instructions, DSM, and TMA. As Table I demonstrates, our work provides comprehensive benchmarks, offering valuable insights for developers seeking to optimize performance and energy efficiency on this latest architecture.
3 Overview of Hopper Architecture
@@
Figure 1: Hopper architecture and new features
In this work, all the above new features will be benchmarked in multiple levels. We select the most representative GPUs of the Ampere, Ada Lovelace, and Hopper architectures, which are A100 PCIe, RTX4090, and H800 PCIe respectively. Their basic hardware properties are shown in Table II. In terms of software configurations, the RTX 4090 utilized driver version 530.30.02 and CUDA version 12.1, while the A100 and H800 employed driver version 560.35.03 and CUDA version 12.6.
TABLE II: Comparison of the Properties of the Ampere, Ada Lovelace and Hopper Devices
@@
Mem. Size 40GB 24GB 80GB
Mem. Type HBM2e GDDR6X HBM2e
Mem. Clock Rate 1215 Mhz 10501 Mhz 1593 Mhz
Mem. Bus 5120-bit 384-bit 5120-bit
Mem. Bandwidth 1555 GB/s 1008 GB/s 2039 GB/s
@@
& Shared Memory
per SM
192KB 128KB 256KB
# of Tensor Core
@@
Distributed
Shared Memory
Tensor Memory
Accelerator
4 Memory Subsystem
Figure 2: Latency clocks of different memory scopes
@@ 4.1 Latency
Traditional P-chase benchmarking. We begin by employing the traditional P-chase microbenchmarking method [38, 39] with random strides [40] to investigate fundamental memory hierarchy characteristics. This includes examining the range of latencies observed and identifying the array size that triggers the latency inflection points. We read data in 64-byte chunks, which corresponds to the size of a generic address. For the L1 cache, L2 cache and global memory, we progressively increase the data volume and capture the inflection points in data access latency.
Fig. 2 shows the access latencies of different memory levels, which are the average values of more than one million accesses. Before the first inflection point, the access latency reflects L1 cache, around 32-33 clocks for all three devices. Between the first and final inflection points, the latency can be attributed to L2 cache. For RTX 4090, L2 cache latency is approximately 284.8 clocks. At the RTX 4090’s third inflection point, the data size roughly equals its L2 cache capacity. As the chain data volume increases, the latency stabilizes around 571 clocks. However, in A100 and H800, we observe two inflection points corresponding to two different latencies for L2 cache. For A100, it is about 202.8 to 408 clocks. In the H800, L2 cache latency is around 264.5 to 502 clocks. In the A100 and H800, the third and fourth points are approximately half of the L2 cache size and the full L2 cache size, respectively. The stable latency after the fourth point, reflecting global memory latency, is 566 clocks for A100 and 656 clocks for H800. Table III concludes the memory access latency at different memory levels. For shared memory, we initialize the P-chase array with uniform 4-byte strides and then access this array using a thread to measure latency. Since shared memory and the L1 cache are physically the same memory component on GPU hardware, we observe similar latency results, approximately 29$\sim$31 cycles. When comparing latency across different memory levels, we observe that, on all three devices, the average latency of the L2 cache is approximately 6.5 times higher than that of the L1 cache, while the global memory latency is about 2.1 times higher than that of the L2 cache, with slight variations across devices.
@@
TABLE III: Latency in clock cycles for different memory scopes using traditional P-chase
@@
Fine-grained P-chase benchmarking. We employ fine-grained P-chase [28] to measure the latency of each memory access, revealing patterns in GPU L2 cache and global memory latency. Our analysis is based on the following assumptions: 1) Two L2 cache partitions have the same capacity. 2) When the size of the data allocated for benchmarking is significantly smaller than half of the total L2 cache capacity, we assume that all memory accesses remain within the cache partition, thus, the result of P-chase does not demonstrate global memory latency. 3) Due to the architectural layout of the GPU, there are differences in the physical proximity between streaming multiprocessors (SMs) and L2 cache partitions. Accessing an L2 partition closer to an SM results in lower latency compared to accessing a farther partition; 4) Any measured latency significantly higher than what can be explained by cache capacity, data size being fully cached, or SM proximity to L2 partitions should be attributed to global memory access due to cache misses. The access sequence of fine-grained P-chase aligns with a uniform 32-byte stride.
On the A100, with an array length of 30MB, four distinct latency groups emerged, corresponding to the four categories in the Table IV. In the H800, an 8MB array exhibited two latency types: L2 cache hit near partition (“near hit”) and L2 cache hit far partition (“far hit”). With a 40MB array, three latency groups were observed: L2 cache near partition hit, L2 cache near partition miss (“near miss”), and L2 cache far partition miss (“far miss”). We apply the K-means clustering algorithm to each device’s four data groups, with cluster center results shown in Table IV. We observe that the A100 and H800, with their dual-partitioned L2 caches, exhibit significantly different behavior compared to the RTX 4090, which has an unpartitioned L2 cache. Additionally, their global memory access latency is considerably higher than that of the RTX 4090.
We also compare the results between Tables III and IV. In the traditional P-chase benchmarking, when the length of the array is less than half the capacity of the L2 cache (between the first and second inflection points), the latencies for A100 and H800 are 202.8 and 264.5 clocks, respectively, aligning with the L2 cache near hit values in Table IV. Furthermore, the global memory latency in Table III approximates (L2 cache near miss + far miss) / 2 from Table IV. This outcome aligns with the probabilistic distribution of data accesses across different L2 cache partitions.
TABLE IV: Latency in clock cycles for different memory scopes using fine-grained P-chase

### v2

@@
Abstract.
This study presents a comprehensive multi-level analysis of the NVIDIA Hopper GPU architecture, focusing on its performance characteristics and novel features. We benchmark Hopper’s memory subsystem, highlighting improvements in the L2 partitioned cache and global memory access compared to Ampere and Ada Lovelace. The evaluation of Hopper’s fourth-generation tensor cores reveals the benefits of FP8 precision and asynchronous wgmma instructions for matrix operations. Additionally, we investigate the performance of DPX instructions for dynamic programming, distributed shared memory (DSM) for inter-SM communication, and the Tensor Memory Accelerator (TMA) for asynchronous data movement. Through multi-level evaluation, we discover that the Hopper architecture demonstrates significant acceleration potential in real-world applications. For instance, the asynchronous programming model supported by TMA achieves a 1.5× speedup in matrix multiplication, FP8 delivers nearly double the performance of FP16, and DPX instructions accelerate a computational biology algorithm by at least 4.75×. Our findings provide actionable insights for optimizing compute-intensive workloads, from AI training to bioinformatics, on Hopper GPUs.
@@ Abstract.
Instruction Benchmark, Tensor Core, PTX, Hopper, DPX, Tensor Memory Accelerator, Distributed Shared Memory, GEMM with TMA, Smith-Waterman Algorithm, Low-Precision Acceleration, Application-Level Benchmarking.
Qiang Wang and Xiaowen Chu are the corresponding authors.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: General and reference Measurementccs: General and reference Evaluationccs: General and reference Performanceccs: Computer systems organization Parallel architecturesfootnotetext: This work is based on and extends ”Benchmarking and Dissecting the Nvidia Hopper GPU Architecture” published in IPDPS 2024. This article expands previous work by adding new tests on L2 partitioned cache, energy consumption of tensor core instructions, comprehensive distributed shared memory evaluations, application-level DPX instruction tests, and benchmarking TMA performance under asynchronous programming on Hopper. The artifact is available on https://github.com/HPMLL/NVIDIA-Hopper-Benchmark.
1. Introduction
@@
The increasing prevalence of GPUs has significantly accelerated a diverse range of computational workloads, from scientific computing to artificial intelligence, particularly with the rise of large language models (LLMs) such as GPT-3, with over 150 billion parameters (Floridi and Chiriatti, 2020). Modern GPU architectures such as Ampere, Ada, and Hopper, incorporate specialized hardware like tensor cores and high-bandwidth memory systems, purpose-built for AI applications. As a result, GPUs have become indispensable components in high-performance computing clusters.
NVIDIA maintains a two-year release cycle for new GPU architectures, ensuring the continuous integration of novel architectural features. However, the scarcity of detailed micro-architectural specifications hinders precise performance analysis, necessitating deeper investigations to understand the impact of these advancements on application performance. Such comprehensive architectural understanding is fundamental for several critical applications: accurate performance and power modeling for energy-efficient computing(Hong and Kim, 2010; Braun et al., 2021; Wang et al., 2019; Wang and Chu, 2020; Mei et al., 2017; Arafa et al., 2020; van Stigt et al., 2022; Arafa et al., 2019), the development of robust GPU simulators that enable cost-effective research and development (Khairy et al., 2020; Bakhoda et al., 2009; Leng et al., 2013), and the effective optimization of applications across diverse domains from AI training to scientific computing (Bakhoda et al., 2009; Ho and Wong, 2017; Yan et al., 2020b). Without detailed microarchitectural insights, developers cannot fully exploit the potential of modern GPUs, leading to suboptimal performance and energy efficiency.
@@
The evolution of tensor cores (TCs) exemplifies this trend. Introduced in the Volta architecture to accelerate deep neural networks using FP16/FP32 operations, TCs have been progressively enhanced. Ampere and Hopper continue to increase support for different precisions, providing more space for research on LLM. Thus, the reliance on assembly analysis and microbenchmarks from older architectures (Ampere and Turing) highlights the need for Hopper-specific TC research.
@@
Beyond enhanced TCs, the Hopper architecture introduces several key innovations. Dynamic Programming X (DPX) instructions accelerate dynamic programming algorithms, which often rely on comparisons (min/max operations) between previously computed results. Distributed shared memory (DSM) enables direct communication between Streaming Multiprocessors (SMs), facilitating loads, stores, and atomic operations across their respective shared memory blocks. Finally, the Tensor Memory Accelerator (TMA) provides an asynchronous data movement mechanism between hardware components. However, the implementation details and performance characteristics of the TMA remain largely unexplored in the existing literature. A comprehensive understanding of these aspects is crucial for developers seeking to optimize compute-intensive workloads, enable accurate performance modeling, and design efficient algorithms that can fully exploit the potential of modern GPUs. Further research is needed to fully understand the interplay of these features and their combined effect on diverse workloads.
@@
This article extends our previous conference paper (Luo et al., 2024), which explored the basic performance of the memory subsystem, a multi-level comparison of fourth-generation tensor cores with previous architectures, and a simple analysis of DPX and Distributed Shared Memory. In this work, we expand both the depth and breadth of our research. In terms of depth, we further analyze the memory subsystem. After various comparisons, we find that some traditional benchmarks are not suitable for Ampere and Hopper. We conduct new tests on the L2 partitioned cache for Ampere and Hopper, overturning previous data and conclusions. Additionally, we include energy consumption tests for wgmma in the tensor core, revealing the power impact of new instructions. We also perform a comprehensive evaluation of distributed shared memory (DSM), assessing DSM interface costs, expanding cluster numbers in latency tests, and testing throughput with different access patterns and thread block scheduling policies. Furthermore, we incorporate application-level testing to further demonstrate the advantages and limitations of DPX instructions. In terms of breadth, with the rise of asynchronous programming, we benchmark the TMA on Hopper to characterize its performance within asynchronous programming paradigms.
@@
$\bullet$ We systematically benchmark the latency and throughput of Hopper GPUs, including the memory subsystem, tensor cores, distributed shared memory (DSM), tensor memory accelerator (TMA), and dynamic programming instructions (DPX). We compare the latency and throughput of these Hopper-specific units with their predecessors in Ampere and Ada Lovelace architectures, highlighting the performance improvements and trade-offs introduced in Hopper.
@@
2. Related Work
@@
The rapid evolution of GPU architectures necessitates continuous investigation into their microarchitectural details, particularly at the instruction level. This understanding is fundamental for accurate performance and power modeling (Hong and Kim, 2010; Braun et al., 2021; Wang et al., 2019; Wang and Chu, 2020; Mei et al., 2017; Arafa et al., 2020; van Stigt et al., 2022; Arafa et al., 2019), the development of robust simulators (Khairy et al., 2020; Bakhoda et al., 2009; Leng et al., 2013), and the effective optimization of applications (Bakhoda et al., 2009; Ho and Wong, 2017; Yan et al., 2020b). Prior research has addressed these needs across different GPU generations, but the introduction of novel features in recent architectures, such as Hopper, demands a renewed focus.
@@
Early efforts to dissect GPU microarchitecture primarily focused on older generations like Fermi, Kepler, and Maxwell (Wong et al., 2010; Mei et al., 2014; Mei and Chu, 2017). These studies, along with investigations into performance and power characteristics (Jia et al., 2018, 2019) and software-level scheduling frameworks(Shekofteh et al., 2020), laid the groundwork for understanding fundamental GPU behavior and informed subsequent investigations into more recent architectures like Volta and Turing. The emergence of tensor cores (TCs) in Volta, specialized hardware units designed to accelerate matrix operations crucial for AI workloads, marked a significant shift in GPU architecture and spurred a new wave of research.
@@
The importance of fused matrix multiplication accumulation (MMA) in AI has driven extensive research into TCs. Initial studies on Volta’s TCs centered on the legacy wmma API and employed benchmarks using established libraries like CUBLAS and CUTLASS (Markidis et al., 2018; Martineau et al., 2019). Further work extended this analysis to Turing, incorporating preliminary assembly code analysis of wmma instructions (Jia et al., 2018, 2019). However, these initial explorations lacked the comprehensive instruction-level microbenchmarks needed to fully characterize TC performance and numerical behavior.
@@
Subsequent research delved deeper into TC analysis on Volta and Turing, utilizing assembly-level (SASS) benchmarking to uncover optimization opportunities for matrix multiplication, particularly in half-precision (Yan et al., 2020a, b; Raihan et al., 2019). The numerical intricacies of TCs, including rounding modes and subnormal number handling for various data types like TF32, BF16, and FP16, were also investigated (Fasi et al., 2021). However, the limitations of the wmma API, particularly its inability to fully leverage new features like sparse matrix multiplication introduced in Ampere and later architectures, became increasingly apparent (Markidis et al., 2018; Fasi et al., 2021).
@@
The introduction of the mma API with Turing addressed some of these limitations and evolved to support sparse matrix multiplication (mma.sp) on Ampere and beyond. Sun et al. (Sun et al., 2023) conducted a thorough analysis of the mma API on Turing and Ampere, exploring TC performance, numerical behavior, and sparse matrix multiplication capabilities. However, the arrival of Hopper introduced further complexity with the addition of wgmma and an expanded set of mma instructions supporting a broader range of precisions. This necessitates new research to understand the performance characteristics and optimal utilization of these new instructions, a gap addressed by our work.
@@
Beyond TC-specific research, previous work has also addressed energy efficiency in GPUs using static (Hong and Kim, 2009, 2010; Braun et al., 2020; Fan et al., 2019) and dynamic (Wang and Chu, 2020; Guerreiro et al., 2018, 2019) analysis techniques, as well as hybrid approaches (Wang et al., 2024). Application-level benchmarks have explored the impact of DVFS on deep learning workloads (Tang et al., 2019; Wang et al., 2020).
Table 1. The Comparison of the studies of GPGPU Microbenchmark. Items in parentheses indicate expansions over (Luo et al., 2024).
Study Instruction-level Memory subsystem Energy consumption Tensor core Application evaluation Architectural features Target architecture
(Jia et al., 2018; Markidis et al., 2018; Martineau et al., 2019) Volta
(Jia et al., 2019; Raihan et al., 2019; Yan et al., 2020a) Turing,Volta
(Fasi et al., 2021; Svedin et al., [n. d.]; Abdelkhalik et al., 2022; Sun et al., 2023) Asynchronous Copy Ampere,Turing
(Luo et al., 2024) DSM, DPX Ampere, Ada, Hopper
This study ✓(L2 partitioned cache) ✓(wgmma) ✓(DSM, TMA, DPX) DSM, TMA, DPX Ampere, Ada, Hopper
Table 1 summarizes the existing microbenchmark studies across different GPU architectures and compares with our work. In contrast to previous research, our study provides a holistic analysis of Hopper, including its novel TCs, DPX instructions, DSM, and TMA. As Table 1 demonstrates, our work provides comprehensive benchmarks, offering valuable insights for developers seeking to optimize performance and energy efficiency on this latest architecture.
3. Overview of Hopper Architecture
@@
Figure 1. Hopper architecture and new features
In this work, all the above new features will be benchmarked in multiple levels. We select the most representative GPUs of the Ampere, Ada Lovelace, and Hopper architectures, which are A100 PCIe, RTX 4090, and H800 PCIe respectively. Their basic hardware properties are shown in Table 2. In terms of software configurations, the RTX 4090 utilized driver version 530.30.02 and CUDA version 12.1, while the A100 and H800 employed driver version 560.35.03 and CUDA version 12.6.
Table 2. Comparison of the Properties of the Ampere, Ada Lovelace and Hopper Devices
@@
Memory Size 40 GB 24 GB 80 GB
Memory Type HBM2e GDDR6X HBM2e
Memory Clock Rate 1215 MHz 10501 MHz 1593 MHz
Memory Bus 5120-bit 384-bit 5120-bit
Memory Bandwidth 1555 GB/s 1008 GB/s 2039 GB/s
@@
& Shared Memory per SM
192 KB 128 KB 256 KB
# of Tensor Cores
@@
Distributed Shared Memory No No Yes
Tensor Memory Accelerator No No Yes
4. Memory Subsystem
Figure 2. Latency clocks of different memory scopes
@@ 4.1. Latency
Traditional P-chase benchmarking. We begin by employing the traditional P-chase microbenchmarking method (Saavedra and Smith, 1995; Saavedra-Barrera, 1992) with random strides (te42kyfo, [n. d.]) to investigate fundamental memory hierarchy characteristics. This includes examining the range of latencies observed and identifying the array size that triggers the latency inflection points. We read data in 64-byte chunks, which corresponds to the size of a generic address. For the L1 cache, L2 cache and global memory, we progressively increase the data volume and capture the inflection points in data access latency. In both Amperehttps://images.nvidia.cn/aem-dam/en-zz/Solutions/data-center/nvidia-ampere-architecture-whitepaper.pdf and Hopperhttps://resources.nvidia.com/en-us-data-center-overview/gtc22-whitepaper-hopper architectures, the L2 cache is divided into two partitions interconnected by high-speed channels. The GPCs are also partitioned in the same way as the L2 cache. Compared to the commonly used uniform stride in previous work(Luo et al., 2024; Wong et al., 2010), random stride better reflects the existence of two L2 cache partitions.
[…]
@@
Table 3. Latency in clock cycles for different memory scopes using traditional P-chase
@@
Fine-grained P-chase benchmarking. We employ fine-grained P-chase (Mei and Chu, 2017) to measure the latency of each memory access, revealing patterns in GPU L2 cache and global memory latency. Our analysis is based on the following assumptions: 1) Two L2 cache partitions have the same capacity. 2) When the size of the data allocated for benchmarking is significantly smaller than half of the total L2 cache capacity, we assume that all memory accesses remain within the cache partition, thus, the result of P-chase does not demonstrate global memory latency. 3) Due to the architectural layout of the GPU, there are differences in the physical proximity between streaming multiprocessors (SMs) and L2 cache partitions. Accessing an L2 partition closer to an SM results in lower latency compared to accessing a farther partition; 4) Any measured latency significantly higher than what can be explained by cache capacity, data size being fully cached, or SM proximity to L2 partitions should be attributed to global memory access due to cache misses. The access sequence of fine-grained P-chase aligns with a uniform 32-byte stride.
On the A100, with an array length of 30MB, four distinct latency groups emerged, corresponding to the four categories in the Table 4. In the H800, an 8MB array exhibited two latency types: L2 cache hit near partition (“near hit”) and L2 cache hit far partition (“far hit”). With a 40MB array, three latency groups were observed: L2 cache near partition hit, L2 cache near partition miss (“near miss”), and L2 cache far partition miss (“far miss”). We apply the K-means clustering algorithm to each device’s four data groups, with cluster center results shown in Table 4. We observe that the A100 and H800, with their dual-partitioned L2 caches, exhibit significantly different behavior compared to the RTX 4090, which has an unpartitioned L2 cache. Additionally, their global memory access latency is considerably higher than that of the RTX 4090.
We also compare the results between Tables 3 and 4. In the traditional P-chase benchmarking, when the length of the array is less than half the capacity of the L2 cache (between the first and second inflection points), the latencies for A100 and H800 are 202.8 and 264.5 clocks, respectively, aligning with the L2 cache near hit values in Table 4. Furthermore, the global memory latency in Table 3 approximates (L2 cache near miss + far miss) / 2 from Table 4. This outcome aligns with the probabilistic distribution of data accesses across different L2 cache partitions.
Table 4. Latency in clock cycles for different memory scopes using fine-grained P-chase